4 g NaOH is dissolved in 500 g water . The concentration of the solution will be? 

Grade:12

2 Answers

Ravleen Kaur
5 years ago
Hi

Conc of solution= Mass of solute/mass of solution
= 4/504= 0.008

Regards
IHE
Khimraj
5 years ago
there are more than one unit to measure concentration of solution.
but mostly used is mol/L or M
mole of NaOH = 4/40 = 1/10
volume of solution = 500/1000 = ½L
so concentration = 0.1/0.5 = 0.2 M
or concentration = 4/500 = 0.008 (unitless)
Hope it clears.